OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <8, 8, -6>; v = <3, -8, -3>; w = <-4, -2, -6>

OpenStudy (e.mccormick):

Know how dot product works?

OpenStudy (e.mccormick):

For \(\vec{u} = \langle u_1, u_2, u_3 \rangle \) and \(\vec{v} = \langle v_1, v_2, v_3\rangle \) \(\vec{u}\cdot\vec{v} = u_1\cdot v_1 + u_2\cdot v_2 + u_3\cdot v_3 \)